OpenStudy (anonymous):

can anybody please explain to me the steps to solve this? 5g + 4 ( -5 +3g ) = 1 - g

OpenStudy (saifoo.khan):

5g +4(-5) +4(3g) = 1 -g

OpenStudy (anonymous):

First, solve 4 ( -5 +3g) getting -20 +12g 5g -20 + 12g = 1 - g 17g - 20 = 1 -g 16g - 20 = 1 16g = 21 g = 21/16

OpenStudy (saifoo.khan):

that is not the answer.

OpenStudy (anonymous):

wait, it's 21/18

OpenStudy (anonymous):

since you're adding g on both sides, not subtracting

OpenStudy (saifoo.khan):

Yes!

OpenStudy (amistre64):

5g + 4 ( -5 +3g ) = 1 - g ; divide by 4 5g/4 -5 +3g = 1/4 - g/4 ; add like terms 17g/4 -5 = 1/4 - g/4 ; swap sides +g/4 +5 ---------------- 18g/4 = 21/4 ; and multiply by 4/18 g = 21/18

OpenStudy (saifoo.khan):

And, 21/18 =7/6

OpenStudy (amistre64):

nah, it equals 42/36 lol

OpenStudy (saifoo.khan):

Right,

OpenStudy (anonymous):

thank you so much!!! ;D